# SCMAlgebras: a Macaulay2 package to check sequential Cohen-Macaulayness

### v2

@@ Abstract.
We introduce the Macaulay2 package SCMAlgebras. It provides functions for computing deficiency modules and filter ideals, in order to check whether a module or an ideal is sequentially Cohen-Macaulay. After the basic algebraic notions and results are recalled, the main features of the package are described through examples.
@@ 1. Introduction
In this paper, we present a new Macaulay2 [10] package, SCMAlgebras, designed to check whether a given module or ideal satisfies the sequentially Cohen-Macaulay property. This package implements key theoretical results, such as Theorem 2.5, which characterizes sequentially Cohen-Macaulay modules in terms of their deficiency modules, and Proposition 2.7 due to Goodarzi [9], which provides an efficient criterion for verifying the sequentially Cohen-Macaulay property for homogeneous ideal. To support these results, the package introduces new computational tools, such as functions for calculating the modules of deficiency (see Definition 2.3) and generating filter ideals (see Definition 2.6), both of which are crucial for testing the sequentially Cohen-Macaulay property of the input objects.
@@ 2. Mathematical background
    for all $0\leq i\leq d$ the $i$th deficiency module $\omega^{i}(M)$ is either zero or an $i$-dimensional Cohen-Macaulay module;
@@
    for all $0\leq i<d$ the $i$th deficiency module $\omega^{i}(M)$ is either zero or an $i$-dimensional Cohen-Macaulay module.
@@ 3. The package
In this section we describe the functions of the package SCMAlgebras by showing some examples of usage and algorithms.
@@
The function deficiencyModule(M,i) uses Definition 2.3 and computes the module $\operatorname{Ext}^{n-i}_{S}(M,S(-n))$. According to Remark 2.4, in our case, the function returns the zero module whenever $i<3$ or $i>4$, and the only non vanishing modules of deficiency are $\omega^{3}(M)$ and $\omega^{4}(M)=\omega(M)$, i.e. the canonical module of $M$. Moreover, the function canonicalModule(M) allows to compute the canonical module $\omega(M)$ directly, as the deficiency module corresponding to dimension.
@@
The function filterIdeals computes the filter ideals of $J_{G}$, according to Definition 2.6. Moreover, the function minDim returns the smallest integer such that the filter ideals differ from $J_{G}$.
@@
minDim(I) Computes the minimum dimension of $I$

### v3

@@ Abstract.
We introduce the Macaulay2 package SCMAlgebras. It provides functions for computing the modules of deficiency and the filter ideals, in order to check whether a module or an ideal is sequentially Cohen-Macaulay. After the basic algebraic notions and results are recalled, the main features of the package are described through examples.
@@ 1. Introduction
In this paper, we present a new Macaulay2 [10] package, SCMAlgebras [14], designed to check whether a given module or ideal satisfies the sequentially Cohen-Macaulay property. This package implements key theoretical results, such as Theorem 2.5, which characterizes sequentially Cohen-Macaulay modules in terms of their deficiency modules, and Proposition 2.7 due to Goodarzi [9], which provides an efficient criterion for verifying the sequentially Cohen-Macaulay property for homogeneous ideal. To support these results, the package introduces new computational tools, such as functions for calculating the modules of deficiency (see Definition 2.3) and generating filter ideals (see Definition 2.6), both of which are crucial for testing the sequentially Cohen-Macaulay property of the input objects.
@@ 2. Mathematical background
    for all $0\leq i\leq d$ the $i$th module of deficiency $\omega^{i}(M)$ is either zero or an $i$-dimensional Cohen-Macaulay module;
@@
    for all $0\leq i<d$ the $i$th module of deficiency $\omega^{i}(M)$ is either zero or an $i$-dimensional Cohen-Macaulay module.
@@ 3. The package
In this section we describe the functions of the package SCMAlgebras [14] by showing some examples of usage and algorithms.
@@
The function deficiencyModule(M,i) uses Definition 2.3 and computes the module $\operatorname{Ext}^{n-i}_{S}(M,S(-n))$. According to Remark 2.4, in our case, the function returns the zero module whenever $i<3$ or $i>4$, and the only non vanishing modules of deficiency are $\omega^{3}(M)$ and $\omega^{4}(M)=\omega(M)$, i.e. the canonical module of $M$. Moreover, the function canonicalModule(M) allows to compute the canonical module $\omega(M)$ directly, as the module of deficiency corresponding to dimension.
@@
The function filterIdeals computes the filter ideals of $J_{G}$, according to Definition 2.6. Moreover, the function minimumDimension returns the smallest integer such that the filter ideals differ from $J_{G}$.
@@
minimumDimension(I) Computes the minimum dimension of $I$